OpenStudy (anonymous):

i am wondering is this correct? (x^3+2x^2-3x^2+2)div(x+1) = x^2-2x+2?

OpenStudy (anonymous):

Yes it is.

OpenStudy (anonymous):

means \[-3x\] not \[-3x^2\] yes?

OpenStudy (anonymous):

is the question correct???????????/

OpenStudy (anonymous):

because i think either way it is wrong

OpenStudy (anonymous):

wrong?

OpenStudy (anonymous):

its correct. all abouve me can flutter off.

OpenStudy (anonymous):

lol @mariaisme

OpenStudy (anonymous):

It's correct!

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

No, it's not correct :(

OpenStudy (anonymous):

\[(x^3+2x^2-3x^2+2)\div(x+1)\] this is how u supposed to write it

OpenStudy (anonymous):

Damn it!! What's wrong with me?! Final answer: CORRECT :D

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

The best way to do it is to use long division!

OpenStudy (anonymous):

correct !

OpenStudy (anonymous):

correct !

OpenStudy (anonymous):

ok by why not subtract first?

OpenStudy (anonymous):

The numerator is\(x^3-x^2+2=(x+1)(x^2-2x+2)\). The (x+1) part will get cancelled with denominator and you'll be left with \(x^2-2x+2\).

OpenStudy (anonymous):

\[x^3+2x^2-3x^2+2=x^2-x^2+2\] yes?

OpenStudy (anonymous):

that is why i thought it was messed up. but it is right!

OpenStudy (anonymous):

thank you guys it been helpful